40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 12563     Entry time: Fri Oct 14 18:33:55 2016     In reply to: 12561     Reply to this: 12566
Author: gautam 
Type: Update 
Category: General 
Subject: AS clipping investigations 

[steve,ericq,gautam]

In the afternoon, we took the heavy door off the OMC chamber as well, such that we could trace the AS beam all the way out to the AP table. 

In summary, we determined the following today:

  1. Beam is centered on SRM, as judged by placing the SOS iris on the tower
     
  2. Beam is a little off on OM1 in yaw, but still >2 beam diameters away from the edge of the steering optic, pitch is pretty good
  3. Beam is okay on OM2 
  4. Beam is okay on OM3 - but beam from OM3 to OM4 is perilously close to clipping on the green steering mirror between these two steering optics (see CAD drawing). We think this is where whatever effect of the SR2 hysteresis shows up first.
  5. Beam is a little low and a little to the left on OM4 (the first PZTJena mirror)
  6. Beam is well clear of other optics in the BS PRM chamber on the way from OM4 to OM5 in the OMC chamber
  7. Beam is a little low and a little to the left of OM5 in the OMC chamber. This is the second PZTJena mirror. We are approximately 1 beam diameter away from clipping on this 1" optic
    Link to IMG_2289.JPG
  8. Beam is off center on OMPO-OMMTSM partially transmissive optic, but because this is a 2" optic, the room for error is much more
    Link to IMG_2294.JPG
  9. Beam is well clear of optics on OMC table on the way from OMPO-OMMTSM to OM6, the final steering mirror bringing the AS beam out onto the table
  10. Beam is low and to the left on OM6. It is pretty bad here, we are < 1 beam diameter away from clipping on this optic, this along with the near miss on the BS/PRM chamber are the two most precarious positions we noticed today, consistent with the hypothesis in this elog that there could be multiple in vacuum clipping points
    Link to IMG_2306.JPG
  11. Beam clears the mirror just before the window pretty confortably (see photo, CAD drawing). But this mirror is not being used for anything useful at the moment. More importantly, there is some reflection off the window back onto this mirror frame which is then scattering and creating some ghost beams, so this could explain the anomalous ASDC behaviour Koji and Yutaro saw. In any case, I would favour removing this mirror since it is serving no purpose at the moment.
    Link to IMG_2310.JPG

Attachment #5 is extracted from the 40m CAD drawing which was last updated in 2012. It shows the beam path for the output beam from the BS all the way to the table (you may need to zoom in to see some labels. The drawing may not be accurate for the OMC chamber but it does show all the relevant optics approximately in their current positions.

EQ will put up photos from the ITMY and BS/PRM chambers.

Plan for Monday: Reconfirm all the findings from today immediately after running the dither alignment so that we can be sure that the ITMs are well-aligned. Then start at OM1 and steer the beam out of the chambers, centering the beam as best as possible given other constraints on all the optics sequentially. All shutters are closed for the weekend, though I left the SOS iris in the chamber...

Here is the link to the Picasa album with a bunch of photos from the OMC chamber prior to us making any changes inside it - there are also some photos in there of the AS beam path inside the OMC chamber...

Attachment 1: IMG_2289.JPG  1.271 MB  Uploaded Sat Oct 15 19:09:09 2016  | Hide | Hide all
IMG_2289.JPG
Attachment 2: IMG_2294.JPG  1.890 MB  Uploaded Sat Oct 15 19:10:21 2016  | Hide | Hide all
IMG_2294.JPG
Attachment 3: IMG_2306.JPG  2.171 MB  Uploaded Sat Oct 15 19:11:20 2016  | Hide | Hide all
IMG_2306.JPG
Attachment 4: IMG_2310.JPG  1.375 MB  Uploaded Sat Oct 15 19:14:34 2016  | Hide | Hide all
IMG_2310.JPG
Attachment 5: ASBeamClipping.pdf  399 kB  Uploaded Sun Oct 16 12:29:11 2016  | Hide | Hide all
ASBeamClipping.pdf
ELOG V3.1.3-